The distribution function of a discrete random variable X is given by

Fx(k) = "\begin{Bmatrix}
 0, & k < 1 \\
 0.1, & 1 \leq k <2 \\
 0.3, & 2 \leq k <3\\
0.7, & 3 \leq k <4 \\
0.8 & 4 \leq k < 5\\
1 & k \geq 5
\end{Bmatrix}"



1) What are the possible values of X?

2) Find the probability mass function of X.



1
Expert's answer
2022-05-25T12:19:39-0400

1) The possible values of "X" are "1,2,3,4,5"

2)


"F(1)=P(X\le1)=p(1)=0.1"

"F(2)=P(X\le2)=p(1)+p(2)=0.3"


"p(2)=0.3-0.1=0.2"

"F(3)=P(X\le3)=p(1)+p(2)+p(3)=0.7"


"p(3)=0.7-0.3=0.4"

"F(4)=P(X\le4)=p(1)+p(2)+p(3)"

"+p(4)=0.8"


"p(4)=0.8-0.7=0.1"

"F(5)=P(X\le5)=p(1)+p(2)+p(3)"

"+p(4)+p(5)=1"


"p(5)=1-0.8=0.2"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 1 & 2 & 3 & 4 & 5 \\ \hline
 f(x) & 0.1 & 0.2 & 0.4 & 0.1 & 0.2 \\

\end{array}"